OpenStudy (anonymous):

I WILL MEDAL** Joshua is the place kicker for his college football team. Last season he kicked 42 times and never missed. Each field goal scored 3 points, and each extra point scored 1 point, for a total of 86 points. How many field goals did Joshua kick? A) 19 B) 20 C) 22 D) 23

OpenStudy (anonymous):

lets call the number of field goals \(x\) then since he kicked \(42\) times the number of extra points is \(42-x\)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

each field goal is worth 3 for a total of \(3x\) and each extra point 1 for a total of \(1\times (42-x)\) or just \(42-x\)

OpenStudy (anonymous):

I did by 42-z

OpenStudy (anonymous):

since you know the total of all points scored is \(86\) you can solve \[3x+42-x=86\]

OpenStudy (anonymous):

oops actually 3 steps combine like terms subtract 42 divide by 2

OpenStudy (anonymous):

So wha i need to find out what is z equal then thats the answer?

OpenStudy (anonymous):

if you used z that is fine equation will be \[3z+42-z=86\]

OpenStudy (anonymous):

*what

OpenStudy (anonymous):

@triciaal do you think you know the answer?

OpenStudy (anonymous):

i should hope so

OpenStudy (anonymous):

Oh ok :)

OpenStudy (anonymous):

can you solve it?

OpenStudy (anonymous):

Ok i can solve it but i just need to know what the combine like terms are then i can do it

OpenStudy (anonymous):

you have \(3z-z\) what is that?

OpenStudy (anonymous):

Ok then i think 42 - y = x 42 - 22 = x 20 = x So, he kicked 22 field goals and 20 extra points.

OpenStudy (anonymous):

@satellite73 right?

OpenStudy (triciaal):

yes